# How to Train Long-Context Language Models (Effectively)

### v3

@@ 6.2 ProLong performance
We first verify that ProLong preserves the base model’s short-context performance in §B.7. We then present the final HELMET evaluation results of ProLong in Table 10. We use all available HELMET tasks here and please refer to Yen et al. (2025) for more details. We compare to a number of frontier long-context LMs, namely MegaBeam [^6] https://huggingface.co/aws-prototyping/MegaBeam-Mistral-7B-512k., Llama-3.1 (Dubey et al., 2024), Qwen2 (Yang et al., 2024a), Phi-3 (Abdin et al., 2024), Mistral-Nemo [^7] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407., Jamba-1.5 (Lenz et al., 2024), Claude-3.5-Sonnet (Anthropic, 2024), Gemini-1.5 (Reid et al., 2024), and GPT-4o (Achiam et al., 2023).
@@
Since most existing models do not support more than 128K tokens, to showcase ProLong’s 512K context length, we stress test ProLong on the QA and summarization tasks from 32K to 512K888 In QA and summarization, we truncate the documents at the evaluation length; hence an effective long-context model should demonstrate better performance on longer lengths.. §6.2 shows that ProLong continues to improve at a longer context window.
@@
Table 12: Results on the NoCha benchmark (Karpinska et al., 2024). [^10] https://github.com/marzenakrp/nocha. NoCha has a private test set and all evaluation is done by the NoCha authors. Hence, we report models from Table 10 that are also on the NoCha leaderboard. ProLong is the only model that achieves above-random performance in the $<$75K category and it consistently beats Llama-3.1. Different from the original NoCha leaderboard, we report the average accuracy over all test instances without filtering the test examples based on the model’s context window lengths.
Model Max Len. $<$75K 75K-127K 127K-180K $>$180K
@@
Besides HELMET, we also evaluate our models on NoCha (Karpinska et al., 2024)—a claim verification dataset on 67 recently published English fictional books. We chose this dataset because (1) it minimizes the data contamination problem as all the books are unlikely to exist in the model pre-training data; (2) all the claims are written by human readers and require global reasoning. Each test instance contains two contradictory claims, and the models must correctly judge both to pass.
footnote 10 demonstrates the NoCha evaluation results. Among 10B-scale models, ProLong achieves the best accuracy on the extremely long test instances ($>$180K); on test instances $<$75K tokens, ProLong significantly outperforms other models and is the only model that is better than random guessing (25%). This further showcases the strength of our training recipe and the ProLong model.
@@ 7 Related Work
Many works explore extending the LM context windows with minimal training, either by position extrapolation (Chen et al., 2023; Peng et al., 2024; Chen et al., 2024; Ding et al., 2024b; Liu et al., 2024a; Zhang et al., 2024b; Zhu et al., 2024; Zhao et al., 2024; Wu et al., 2024; Hu et al., 2024b) or manipulating the attention patterns (Chen et al., 2024; Xiao et al., 2024b; a; Bertsch et al., 2023; Jin et al., 2024). Yoshida et al. (2020); Choromanski et al. (2021); Chevalier et al. (2023) instead explore the idea of compressing the long contexts into shorter forms. However, Fu et al. (2024); Lu et al. (2024) show that using full attention, applying simple position extrapolation, and fine-tuning the model on long documents reach much stronger results.
Llama 3.1 (Dubey et al., 2024) and Jamba (Lieber et al., 2024) achieve long-context capabilities by adding a long-context continued training stage between standard pre-training and supervised fine-tuning, which is the setting we follow. Fu et al. (2024) study the data engineering for this setting and argue that 0.5B tokens of domain-balanced, length-upsampled data is sufficient for acquiring the long-context recall ability—which we show is not sufficient if a more holistic evaluation is taken. Xiong et al. (2023); Dubey et al. (2024); Lieber et al. (2024); Xiong et al. (2024); An et al. (2024b); Bai et al. (2024a) also adopt synthetically-generated long data in the SFT stage; however, we find that using standard, short-context instruction data achieves the best long-context results in our setting.
@@
There have been many efforts in designing more efficient architectures, for example, linear attention/RNNs (Gu & Dao, 2023; Dao & Gu, 2024; Ma et al., 2022; Sun et al., 2023; Peng et al., 2023; Yang et al., 2024b), and alternative attention architectures (Rubin & Berant, 2023; Sun et al., 2024; Yen et al., 2024). However, they often require training from scratch and many have the inherent limitations in terms of long-context recall (Jelassi et al., 2024; Arora et al., 2024). Recent works explore hybrid models (Waleffe et al., 2024; Lieber et al., 2024) or distilling existing LMs into hybrid models (Wang et al., 2024) and show promising results.
@@
Many benchmarks have been proposed for long-context evaluation (Shaham et al., 2023; Hsieh et al., 2024; Krishna et al., 2023; Zhang et al., 2024a; An et al., 2024a; Bai et al., 2024b) There are works studying particular aspects of long-context LMs as well, such as positional bias (Liu et al., 2024b), in-context learning (Bertsch et al., 2024; Li et al., 2024b), and book-length summarization (Kim et al., 2024). In this work, we follow Yen et al. (2025) for its diverse application coverage and reliable evaluations.
@@ Acknowledgments
We acknowledge Mengzhou Xia, Zexuan Zhong, Samyak Gupta, Dan Friedman, Yihe Dong, Abhishek Panigrah, Adithya Bhaskar, Colin Wang, Carlos Jimenez, and other members of Princeton Language and Intelligence for their helpful feedback and discussion. We also thank Luca Soldaini for providing comments on a draft. We thank Marzena Karpinska, Tanya Goyal, and Mohit Iyyer for their help with the NoCha evaluation. Tianyu Gao is supported by an IBM PhD Fellowship. This work is gratefully supported by an NSF CAREER award (IIS-2239290), a grant from Intel, and Microsoft Azure credits through the “Accelerate Foundation Models Academic Research” Initiative.
@@ A.1 Evaluation
Table 13: The details for our long-context evaluation following HELMET (Yen et al., 2025).
@@
Recall SubEM Given a randomly-generated long JSON file and a key, retrieve the corresponding value (Liu et al., 2024b).
RAG SubEM Given a question and many retrieved Wikipedia documents (shuffled), answer the question (Liu et al., 2024b). Datasets: NaturalQuestion (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and PopQA (Mallen et al., 2023).
Re-rank nDCG@10 Given a query and many retrieved documents (shuffled), re-rank the top-10 documents. Datasets: MSMARCO (Bajaj et al., 2016).
ICL Accuracy Datasets selected from Bertsch et al. (2024): TREC coarse, TREC fine (Hovy et al., 2001), NLU (Liu et al., 2021), Banking77 (Casanueva et al., 2020), and Clinc-150 (Larson et al., 2019).
QA GPT-4o score Given a book, answer the question. Datasets (# tokens): NarrativeQA (medium: 73K; max: 518K; Kočiský et al., 2018).
Summ. GPT-4o score Summarize a given legal document. Datasets (# tokens): Multi-LexSum (medium: 90K; max: 5M; Shen et al., 2022)
Table 13 shows all the datasets we used for the long-context evaluation from HELMET (Yen et al., 2025). Note that we did not use all the datasets from HELMET for efficiency reasons and we also do not want to overfit to HELMET. We highlight some of the evaluation protocol improvements that HELMET implemented compared to previous benchmarks here:
@@
  • Better synthetic recall tasks. As shown in HELMET, needle-in-a-haystack (Kamradt, 2024) is mostly saturated because of its simplicity—the model only needs to find a needle in some irrelevant context. We instead use the more challenging JSON KV task, first proposed in Liu et al. (2024b) and included in HELMET, where the model is required to find the corresponding value to a given key among a large JSON file.
  • Using class-balanced demonstrations and abstract labels for ICL. To disentangle models’ ability of learning from demonstrations from their pre-training bias of the task or the dataset label distribution (Pan et al., 2023), HELMET samples the same number of demonstrations for each class and uses number labels (1, 2, …) instead of natural-language labels (e.g., location, description, …).
  • Model-based evaluation for long-context QA and summarization. Instead of using traditional metrics like ROUGE (which has shown to be poorly indicative of the real model performance: Deutsch & Roth, 2021; Deutsch et al., 2022; Goyal et al., 2023; Chang et al., 2024), HELMET uses model-based evaluations to compare the reference answer and the model output. For QA, HELMET uses GPT-4o to score the model output given the question and the reference answer at a 0-3 scale. For summarization, HELMET takes a similar approach as Zhang & Bansal (2021); Gao et al. (2023): it first uses GPT-4o to decompose the reference summary into atomic claims; then it uses GPT-4o to check whether each reference atomic claim is covered by the model output (recall) and whether each sentence in the model output is covered by the reference summary (precision). Yen et al. (2025) show that the model-based evaluation correlates with human perceptions significantly better than traditional metrics.
@@ A.2 Data processing
We list all the data sources we have explored in our ablations and main experiments here: the Stack (Kocetkov et al., 2023), SlimPajama (Together, 2023; Soboleva et al., 2023), FineWeb (we use the 2023-50 snapshot), FineWeb-Edu (we use a random sample) (Penedo et al., 2024), Tulu-v2 (Ivison et al., 2023), OpenWebMath (Paster et al., 2024), textbooks (Chevalier et al., 2024), and Dolma (Soldaini et al., 2024). The Books, StackExchange, and ArXiv data are from SlimPajama. The Wikipedia data are from Dolma.
@@
For 512K length, we use a mix of 64K and 512K long data. For the ratio of 64K/512K data, we choose 50%/50% for code and 83%/17%, which are roughly chosen according to the natural availability of very long data, i.e., there are relatively fewer books of length 512K than code repositories. One benefit of retaining 64K-long documents is that we can process these without sequence parallelism and the associated communication overhead. We use a slightly different long data mixture in our ablations (Table 5) and our main ProLong experiment (Table 9). For the final model, we mix 3% textbooks into the long-context training data. The textbooks are open-source resources from libretexts.org, collected and made available by Chevalier et al. (2024). We pre-process the data by concatenating chapters from the same text books, as well as books from the same subject areas. This results in extremely long sequences which we pack into contexts of either 64K or 512K tokens. Though we do not have an ablation for adding this data due to limited resources, we believe that it should have a slight positive effect to the final model performance as textbooks are highly educational long-context data.
@@ A.3 Implementation details
We use various open-source packages and tools for the ProLong training and evaluation. We use PyTorch (Paszke et al., 2019) and Hugging Face transformers (Wolf et al., 2020) for the model training. We use mosaic-streaming (Mosaic ML, 2022) for loading and mixing the data and FlashAttention 2 (Dao, 2024) for efficient attention implementation. We implement sequence parallelism based on DeepSpeed-Ulysses (Jacobs et al., 2023) across groups of 8 GPUs on the same node. We only perform distributed attention if it is necessary, i.e., only on sequences of 512K length. For long-context evaluation, we use HELMET (Yen et al., 2025) and for short-context evaluation, we use lm-eval-harness (Gao et al., 2021).
@@
Since we do document masking in attention (§6), we use the variable-length attention implementation from FlashAttention 2 (Dao, 2024) to speed up long-context training: for sequences that are concatenations of multiple short documents, instead of computing the full attention with masking, we instead compute the attention for each individual document. Since the complexity of attention is quadratic to the sequence length, this improves the training speed. However, the improvement is negligible in a distributed training setting with FSDP, since GPUs processing short sequence batches have to wait on other GPUs processing long sequences. We therefore implement a smart batching algorithm: In our setting, a gradient step usually consists of multiple gradient accumulation steps, where each device processes a smaller minibatch. We sort all the minibatches per training step by the sum of the squared lengths of documents in the sequence. This leads to more balanced sequence lengths across the GPUs and effective speedups, as can be seen in Table 15, without affecting the gradient updates or loss during training. However, the efficiency gains are diminished when training with more GPUs, as this reduces the number of gradient accumulation steps.
@@
We found that in the SFT stage, the distribution of the training tokens (in SFT, the tokens from the instructions are masked out and the models are only trained on the responses) on each GPU device can be extremely imbalanced, especially when there is synthetic data (most tokens in a synthetic data instance are from the instruction). Conventional all-reduce loss in distributed training averages over the sequences instead of valid tokens, which skews the optimization and also our control over the domain proportions. Instead, we change the all-reduce loss to be the average over all valid training tokens. Bai et al. (2024a) implements their SFT loss in a similar way.
@@ B.1 Position extrapolation
Xiong et al. (2023); emozilla (2023) show that changing the RoPE frequency base to a larger value in continual long-context pre-training or in inference time can improve the long-context performance. emozilla (2023) suggests that one should scale the frequency base by a factor of $t^{\frac{d}{d-2}}$, where $t$ is the ratio between the target sequence length and the original LM length, and $d$ is the attention head dimension.
We conduct ablation studies, at both 64K (same as our standard ablation setting as specified in §A.4) and 512K (starting from ProLong-64K and training with the 512K data mixture for 5B tokens) sequence lengths, on what frequency bases we should use. §B.1 and §B.1 show the results. We first see that using the original 500,000 frequency base from Llama-3 leads to significant performance degradation. While dynamic NTK suggests $4\times 10^{6}$, we find that further scaling it to $8\times 10^{6}$ leads to better performance. Similar, we see that when scaling the 64K model to 512K, while dynamic NTK suggests a $64\times 10^{6}$ frequency base, much larger frequency bases ($128\times 10^{6}$ and $256\times 10^{6}$) lead to better performance. We use $8\times 10^{6}$ for 64K and $128\times 10^{6}$ for 512K for our final ProLong models.
Table 18: Ablation study on RoPE frequency base at a maximum training length of 64K. Dynamic NTK (emozilla, 2023) roughly suggests to use 4m as the frequency base.
RoPE Base (×$\times 10^{6}$) Long-Context Short-Context
@@
Table 19: Ablation study on RoPE frequency base at a maximum training length of 512K. Dynamic NTK (emozilla, 2023) roughly suggests to use $64\times 10^{6}$ as the frequency base.
RoPE Base (×$\times 10^{6}$) Long-Context Short-Context

### v4

@@ 6.2 ProLong performance
We first verify that ProLong preserves the base model’s short-context performance in §B.7. We then present the final HELMET evaluation results of ProLong in Table 10. We use all available HELMET tasks here and please refer to yen2024helmet for more details. We compare to a number of frontier long-context LMs, namely MegaBeam [^6] https://huggingface.co/aws-prototyping/MegaBeam-Mistral-7B-512k., Llama-3.1 (dubey2024llama), Qwen2 (yang2024qwen2), Phi-3 (abdin2024phi), Mistral-Nemo [^7] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407., Jamba-1.5 (team2024jamba15), Claude-3.5-Sonnet (anthropic2024claude), Gemini-1.5 (reid2024gemini), and GPT-4o (achiam2023gpt).
@@
Since most existing models do not support more than 128K tokens, to showcase ProLong’s 512K context length, we stress test ProLong on the QA and summarization tasks from 32K to 512K [^8] In QA and summarization, we truncate the documents at the evaluation length; hence an effective long-context model should demonstrate better performance on longer lengths.. Table 11 shows that ProLong continues to improve at a longer context window.
@@
Table 12: Results on the NoCha benchmark (karpinska2024one). [^10] https://github.com/marzenakrp/nocha. NoCha has a private test set and all evaluation is done by the NoCha authors. Hence, we report models from Table 10 that are also on the NoCha leaderboard. ProLong is the only model that achieves above-random performance in the <<75K category and it consistently beats Llama-3.1. Different from the original NoCha leaderboard, we report the average accuracy over all test instances without filtering the test examples based on the model’s context window lengths.
Model Max Len. <<75K 75K-127K 127K-180K >>180K
@@
Besides HELMET, we also evaluate our models on NoCha (karpinska2024one)—a claim verification dataset on 67 recently published English fictional books. We chose this dataset because (1) it minimizes the data contamination problem as all the books are unlikely to exist in the model pre-training data; (2) all the claims are written by human readers and require global reasoning. Each test instance contains two contradictory claims, and the models must correctly judge both to pass.
footnote 10 demonstrates the NoCha evaluation results. Among 10B-scale models, ProLong achieves the best accuracy on the extremely long test instances (>>180K); on test instances <<75K tokens, ProLong significantly outperforms other models and is the only model that is better than random guessing (25%). This further showcases the strength of our training recipe and the ProLong model.
@@ 7 Related Work
Many works explore extending the LM context windows with minimal training, either by position extrapolation (chen2023extending; peng2024yarn; chen2024longlora; ding2024longrope; liu20242; zhang2024extending; zhu2024pose; zhao2024longskywork; wu2024long; hu2024longrecipe) or manipulating the attention patterns (chen2024longlora; xiao2024efficient; xiao2024infllm; bertsch2023unlimiformer; jin2024llm). yoshida2020adding; choromanski2021rethinking; chevalier-etal-2023-adapting instead explore the idea of compressing the long contexts into shorter forms. However, fu2024data; lu2024controlled show that using full attention, applying simple position extrapolation, and fine-tuning the model on long documents reach much stronger results.
Llama 3.1 (dubey2024llama) and Jamba (lieber2024jamba) achieve long-context capabilities by adding a long-context continued training stage between standard pre-training and supervised fine-tuning, which is the setting we follow. fu2024data study the data engineering for this setting and argue that 0.5B tokens of domain-balanced, length-upsampled data is sufficient for acquiring the long-context recall ability—which we show is not sufficient if a more holistic evaluation is taken. xiong2023effective; dubey2024llama; lieber2024jamba; xiong2024artificial; an2024make; bai2024longalign also adopt synthetically-generated long data in the SFT stage; however, we find that using standard, short-context instruction data achieves the best long-context results in our setting.
@@
There have been many efforts in designing more efficient architectures, for example, linear attention/RNNs (gu2023mamba; dao2024transformers; ma2022mega; sun2023retentive; peng-etal-2023-rwkv; yang2024gated), and alternative attention architectures (rubin2023longrange; sun2024you; yen-etal-2024-long). However, they often require training from scratch and many have the inherent limitations in terms of long-context recall (jelassi2024repeat; arora2024simple). Recent works explore hybrid models (waleffe2024empirical; lieber2024jamba) or distilling existing LMs into hybrid models (wang2024mamba) and show promising results.
@@
Many benchmarks have been proposed for long-context evaluation (shaham-etal-2023-zeroscrolls; hsieh2024ruler; krishna-etal-2023-longeval; zhang-etal-2024-bench; an-etal-2024-l; bai-etal-2024-longbench) There are works studying particular aspects of long-context LMs as well, such as positional bias (Liu2023LostIT), in-context learning (bertsch2024context; li2024long), and book-length summarization (kim2024fables). In this work, we follow yen2024helmet for its diverse application coverage and reliable evaluations.
@@ Acknowledgments
We acknowledge Mengzhou Xia, Zexuan Zhong, Samyak Gupta, Dan Friedman, Yihe Dong, Abhishek Panigrah, Adithya Bhaskar, Colin Wang, Carlos Jimenez, and other members of Princeton Language and Intelligence for their helpful feedback and discussion. We also thank Luca Soldaini for providing comments on a draft. We thank Marzena Karpinska, Tanya Goyal, and Mohit Iyyer for their help with the NoCha evaluation. Tianyu Gao is supported by an IBM PhD Fellowship. This work is gratefully supported by an NSF CAREER award (IIS-2239290), a grant from Intel, Cisco Research, and Microsoft Azure credits through the “Accelerate Foundation Models Academic Research” Initiative.
@@ A.1 Evaluation
Table 13: The details for our long-context evaluation following HELMET (yen2024helmet).
@@
Recall SubEM Given a randomly-generated long JSON file and a key, retrieve the corresponding value (Liu2023LostIT).
RAG SubEM Given a question and many retrieved Wikipedia documents (shuffled), answer the question (Liu2023LostIT). Datasets: NaturalQuestion (kwiatkowski2019natural), HotpotQA (yang-etal-2018-hotpotqa), and PopQA (mallen-etal-2023-trust).
Re-rank nDCG@10 Given a query and many retrieved documents (shuffled), re-rank the top-10 documents. Datasets: MSMARCO (bajaj2016ms).
ICL Accuracy Datasets selected from bertsch2024context: TREC coarse, TREC fine (hovy-etal-2001-toward), NLU (liu2021benchmarking), Banking77 (casanueva-etal-2020-efficient), and Clinc-150 (larson-etal-2019-evaluation).
QA GPT-4o score Given a book, answer the question. Datasets (# tokens): NarrativeQA (medium: 73K; max: 518K; kocisky2018narrativeqa).
Summ. GPT-4o score Summarize a given legal document. Datasets (# tokens): Multi-LexSum (medium: 90K; max: 5M; multilexsum)
Table 13 shows all the datasets we used for the long-context evaluation from HELMET (yen2024helmet). Note that we did not use all the datasets from HELMET for efficiency reasons and we also do not want to overfit to HELMET. We highlight some of the evaluation protocol improvements that HELMET implemented compared to previous benchmarks here:
@@
  • Better synthetic recall tasks. As shown in HELMET, needle-in-a-haystack (gkamradt_llmtest_needleinahaystack_2024) is mostly saturated because of its simplicity—the model only needs to find a needle in some irrelevant context. We instead use the more challenging JSON KV task, first proposed in Liu2023LostIT and included in HELMET, where the model is required to find the corresponding value to a given key among a large JSON file.
  • Using class-balanced demonstrations and abstract labels for ICL. To disentangle models’ ability of learning from demonstrations from their pre-training bias of the task or the dataset label distribution (pan-etal-2023-context), HELMET samples the same number of demonstrations for each class and uses number labels (1, 2, …) instead of natural-language labels (e.g., location, description, …).
  • Model-based evaluation for long-context QA and summarization. Instead of using traditional metrics like ROUGE (which has shown to be poorly indicative of the real model performance: deutsch-roth-2021-understanding; deutsch-etal-2022-examining; goyal2023newssummarizationevaluationera; chang2024booookscore), HELMET uses model-based evaluations to compare the reference answer and the model output. For QA, HELMET uses GPT-4o to score the model output given the question and the reference answer at a 0-3 scale. For summarization, HELMET takes a similar approach as zhang-bansal-2021-finding; gao2023alce: it first uses GPT-4o to decompose the reference summary into atomic claims; then it uses GPT-4o to check whether each reference atomic claim is covered by the model output (recall) and whether each sentence in the model output is covered by the reference summary (precision). yen2024helmet show that the model-based evaluation correlates with human perceptions significantly better than traditional metrics.
@@ A.2 Data processing
We list all the data sources we have explored in our ablations and main experiments here: the Stack (kocetkov2023the), SlimPajama (together2023redpajama; cerebras2023slimpajama), FineWeb (we use the 2023-50 snapshot), FineWeb-Edu (we use a random sample) (penedo2024finewebdatasetsdecantingweb), Tulu-v2 (ivison2023camels), OpenWebMath (paster2024openwebmath), textbooks (chevalier2024language), and Dolma (soldaini-etal-2024-dolma). The Books, StackExchange, and ArXiv data are from SlimPajama. The Wikipedia data are from Dolma.
@@
For 512K length, we use a mix of 64K and 512K long data. For the ratio of 64K/512K data, we choose 50%/50% for code and 83%/17%, which are roughly chosen according to the natural availability of very long data, i.e., there are relatively fewer books of length 512K than code repositories. One benefit of retaining 64K-long documents is that we can process these without sequence parallelism and the associated communication overhead. We use a slightly different long data mixture in our ablations (Table 5) and our main ProLong experiment (Table 9). For the final model, we mix 3% textbooks into the long-context training data. The textbooks are open-source resources from libretexts.org, collected and made available by chevalier2024language. We pre-process the data by concatenating chapters from the same text books, as well as books from the same subject areas. This results in extremely long sequences which we pack into contexts of either 64K or 512K tokens. Though we do not have an ablation for adding this data due to limited resources, we believe that it should have a slight positive effect to the final model performance as textbooks are highly educational long-context data.
@@ A.3 Implementation details
We use various open-source packages and tools for the ProLong training and evaluation. We use PyTorch (paszke2019pytorch) and Hugging Face transformers (wolf2020transformers) for the model training. We use mosaic-streaming (mosaicml2022streaming) for loading and mixing the data and FlashAttention 2 (dao2024flashattention) for efficient attention implementation. We implement sequence parallelism based on DeepSpeed-Ulysses (jacobs2023deepspeed) across groups of 8 GPUs on the same node. We only perform distributed attention if it is necessary, i.e., only on sequences of 512K length. For long-context evaluation, we use HELMET (yen2024helmet) and for short-context evaluation, we use lm-eval-harness (eval-harness).
@@
Since we do document masking in attention (§6), we use the variable-length attention implementation from FlashAttention 2 (dao2024flashattention) to speed up long-context training: for sequences that are concatenations of multiple short documents, instead of computing the full attention with masking, we instead compute the attention for each individual document. Since the complexity of attention is quadratic to the sequence length, this improves the training speed. However, the improvement is negligible in a distributed training setting with FSDP, since GPUs processing short sequence batches have to wait on other GPUs processing long sequences. We therefore implement a smart batching algorithm: In our setting, a gradient step usually consists of multiple gradient accumulation steps, where each device processes a smaller minibatch. We sort all the minibatches per training step by the sum of the squared lengths of documents in the sequence. This leads to more balanced sequence lengths across the GPUs and effective speedups, as can be seen in Table 15, without affecting the gradient updates or loss during training. However, the efficiency gains are diminished when training with more GPUs, as this reduces the number of gradient accumulation steps.
@@
We found that in the SFT stage, the distribution of the training tokens (in SFT, the tokens from the instructions are masked out and the models are only trained on the responses) on each GPU device can be extremely imbalanced, especially when there is synthetic data (most tokens in a synthetic data instance are from the instruction). Conventional all-reduce loss in distributed training averages over the sequences instead of valid tokens, which skews the optimization and also our control over the domain proportions. Instead, we change the all-reduce loss to be the average over all valid training tokens. bai2024longalign implements their SFT loss in a similar way.
@@ B.1 Position extrapolation
xiong2023effective; dynamicntk show that changing the RoPE frequency base to a larger value in continual long-context pre-training or in inference time can improve the long-context performance. dynamicntk suggests that one should scale the frequency base by a factor of $t^{\frac{d}{d-2}}$, where $t$ is the ratio between the target sequence length and the original LM length, and $d$ is the attention head dimension.
We conduct ablation studies, at both 64K (same as our standard ablation setting as specified in §A.4) and 512K (starting from ProLong-64K and training with the 512K data mixture for 5B tokens) sequence lengths, on what frequency bases we should use. Table 18 and Table 19 show the results. We first see that using the original 500,000 frequency base from Llama-3 leads to significant performance degradation. While dynamic NTK suggests $4\times 10^{6}$, we find that further scaling it to $8\times 10^{6}$ leads to better performance. Similar, we see that when scaling the 64K model to 512K, while dynamic NTK suggests a $64\times 10^{6}$ frequency base, much larger frequency bases ($128\times 10^{6}$ and $256\times 10^{6}$) lead to better performance. We use $8\times 10^{6}$ for 64K and $128\times 10^{6}$ for 512K for our final ProLong models.
Table 18: Ablation study on RoPE frequency base at a maximum training length of 64K. Dynamic NTK (dynamicntk) roughly suggests to use 4m as the frequency base.
RoPE Base ($\times 10^{6}$) Long-Context Short-Context
@@
Table 19: Ablation study on RoPE frequency base at a maximum training length of 512K. Dynamic NTK (dynamicntk) roughly suggests to use $64\times 10^{6}$ as the frequency base.
RoPE Base ($\times 10^{6}$) Long-Context Short-Context